# 2D Gaussian Splatting for Geometrically Accurate Radiance Fields

### v1

@@ 6. Experiments
We now present an extensive evaluation of our 2D Gaussian Splatting reconstruction method, including appearance and geometry comparison with previous state-of-the-art implicit and explicit approaches. We then analyze the contribution of the proposed components.
@@ 6.1. Implementation
We implement our 2D Gaussian Splatting with custom CUDA kernels, building upon the framework of 3DGS (Kerbl et al., 2023). We extend the renderer to output depth distortion maps, depth maps and normal maps for regularizations. During training, we increase the number of 2D Gaussian primitives following the adaptive control strategy in 3DGS. Since our method does not directly rely on the gradient of the projected 2D center, we hence project the gradient of 3D center $\mathbf{p}_{k}$ onto the screen space as an approximation. Similarly, we employ a gradient threshold of $0.0002$ and remove splats with opacity lower than $0.05$ every $3000$ step. We conduct all the experiments on a single GTX RTX3090 GPU.
Mesh Extraction: To extract meshes from reconstructed 2D splats, we render depth maps of the training views using the median depth value of the splats projected to the pixels and utilize truncated signed distance fusion (TSDF) to fuse the reconstruction depth maps, using Open3D (Zhou et al., 2018). We set the voxel size to $0.004$ and the truncated threshold to 0.02 during TSDF fusion. We also extend the original 3DGS to render depth and employ the same technique for surface reconstruction for a fair comparison.
@@
Courthouse 0.17 0.12 0.28 0.08 0.09 0.13
Ignatius 0.83 0.72 0.89 0.33 0.04 0.44
Meetingroom 0.24 0.20 0.32 0.15 0.01 0.16
Truck 0.45 0.45 0.48 0.26 0.19 0.26
Mean 0.38 0.35 0.50 0.19 0.09 0.30
Time ¿24h ¿24h ¿24h ¿1h 14.3 m 34.2  m
6.2. Comparison
Dataset: We evaluate the performance of our method on various datasets, including DTU (Jensen et al., 2014), Tanks and Temples (Knapitsch et al., 2017), and Mip-NeRF360 (Barron et al., 2022a). The DTU dataset comprises 15 scenes, each with 49 or 69 images of resolution $1600\times 1200$. We use Colmap (Schönberger and Frahm, 2016) to generate a sparse point cloud for each scene and down-sample the images into a resolution of $800\times 600$ for efficiency. We use the same training process for 3DGS (Kerbl et al., 2023) and SuGar (Guédon and Lepetit, 2023) for a fair comparison.
Geometry Reconstruction: In  Table 1 and  Table 3, we compare our geometry reconstruction to SOTA implicit (i.e., NeRF (Mildenhall et al., 2020), VolSDF (Yariv et al., 2021), and NeuS (Wang et al., 2021)), explicit (i.e., 3DGS (Kerbl et al., 2023) and concurrent work SuGaR (Guédon and Lepetit, 2023)) on Chamfer distance and training time using the DTU dataset. Our method outperforms all compared methods in terms of Chamfer distance. Moreover, as shown in Table 2, 2DGS achieves competitive results with SDF models (i.e., NeuS (Wang et al., 2021) and Geo-Neus (Fu et al., 2022)) on the TnT dataset, and significantly better reconstruction than explicit reconstruction methods (i.e., 3DGS and SuGaR). Notably, our model demonstrates exceptional efficiency, offering a reconstruction speed that is approximately 100 times faster compared to implicit reconstruction methods and more than 3 times faster than the concurrent work SuGaR. Our approach can also achieve qualitatively better reconstructions with more appearance and geometry details and fewer outliers, as shown in  Figure 5. We include the geometry reconstruction results in supplementary  Figure 10. Moreover, SDF-based reconstruction methods require predefining the spherical size for initialization, which plays a critical role in the success of SDF reconstruction. Bu contrast, our method leverages radiance field based geometry modeling and is less sensitive to initialization.
Table 3. Performance comparison between 2DGS (ours), 3DGS and SuGaR on the DTU dataset (Jensen et al., 2014). We report the averaged chamfer distance, PSNR, reconstruction time, and model size.
@@ 6.2. Comparison
Appearance Reconstruction: Our method represents 3D scenes as radiance fields, providing high-quality novel view synthesis. In this section, we compare our novel view renderings using the Mip-NeRF360 dataset against baseline approaches, as shown in  Table 4 and  Figure 4. Note that, since the ground truth geometry is not available in the Mip-NeRF360 dataset and we hence focus on quantitative comparison. Remarkably, our method consistently achieves competitive NVS results across state-of-the-art techniques while providing geometrically accurate surface reconstruction.
@@
3DGS 24.24 0.705 0.283 30.99 0.926 0.199
SuGaR 22.76 0.631 0.349 29.44 0.911 0.216
2DGS (Ours) 24.33 0.709 0.284 30.39 0.924 0.182
6.3. Ablations
[…]
Regularization: We first examine the effects of the proposed normal consistency and depth distortion regularization terms. Our model (Table 5 E) provides the best performance when applying both regularization terms. We observe that disabling the normal consistency (Table 5 A) can lead to incorrect orientation, as shown in Figure 6 A. Additionally, the absence of depth distortion (Table 5 B) results in a noisy surface, as shown in Figure 6 B.
@@ 6.3. Ablations
Accuracy $\downarrow$ Completion  $\downarrow$ Averagy $\downarrow$
@@
Mesh Extraction: We now analyze our choice for mesh extraction. Our full model (Table 5 E) utilizes TSDF fusion for mesh extraction with median depth. One alternative option is to use the expected depth instead of the median depth. However, it yields worse reconstructions as it is more sensitive to outliers, as shown in Table 5 C. Further, our approach surpasses screened Poisson surface reconstruction (SPSR)(Table 5 D) (Kazhdan and Hoppe, 2013) using 2D Gaussians’ center and normal as inputs, due to SPSR’s inability to incorporate the opacity and the size of 2D Gaussian primitives.
@@ 7. Conclusion
Limitations: While our method successfully delivers accurate appearance and geometry reconstruction for a wide range of objects and scenes, we also discuss its limitations: First, we assume surfaces with full opacity and extract meshes from multi-view depth maps. This can pose challenges in accurately handling semi-transparent surfaces, such as glass, due to their complex light transmission properties. Secondly, our current densification strategy favors texture-rich over geometry-rich areas, occasionally leading to less accurate representations of fine geometric structures. A more effective densification strategy could mitigate this issue. Finally, our regularization often involves a trade-off between image quality and geometry, and can potentially lead to oversmoothing in certain regions.
Acknowledgement: BH and SG are supported by NSFC #62172279, #61932020, Program of Shanghai Academic Research Leader. ZY, AC and AG are supported by the ERC Starting Grant LEGO-3D (850533) and DFG EXC number 2064/1 - project number 390727645.
@@ Appendix A Details of Depth Distortion
(17) $\small\begin{split}\mathcal{L}&=\sum_{i=0}^{N-1}\sum_{j=0}^{i-1}\omega_{i}\omega_{j}(m_{i}-m_{j})^{2}\\ &=\sum_{i=0}^{N-1}\omega_{i}\left(m_{i}^{2}\sum_{j=0}^{i-1}\omega_{j}+\sum_{j=0}^{i-1}\omega_{j}m_{j}^{2}-2m_{i}\sum_{j=0}^{i-1}\omega_{j}m_{j}\right)\\ &=\sum_{i=0}^{N-1}\omega_{i}\left(m_{i}^{2}A_{i-1}+D^{2}_{i-1}-2m_{i}D_{i-1}\right),\end{split}$
@@ Appendix B Depth calculations
Mean depth:There are two optional depth computations used for our meshing process. The mean (expected) depth is calculated by weighting the intersected depth:
@@
Median depth:We compute the median depth as the largest “visible” depth, considering $T_{i}=0.5$ as the pivot for surface and free space:
@@
Appendix C Additional Results
Our 2D Gaussian Splatting method achieves comparable performance even without the need for regularizations, as Table 7 shows. Additionally, we have provided a breakdown of various per-scene metrics for the MipNeRF360 dataset (Barron et al., 2022b) in Tables 6. Figure 8 presents a comparison of our rendered depth maps with those from 3DGS and MipNeRF360.
@@ Appendix C Additional Results
3DGS 24.71 21.09 26.63 26.45 22.33 31.50 29.07 31.13 32.26 27.24
SuGaR 23.12 19.25 25.43 24.69 21.33 30.12 27.57 29.48 30.59 25.73
Ours 24.82 20.99 26.91 26.41 22.52 30.86 28.45 30.62 31.64 27.03
3DGS 0.729 0.571 0.834 0.762 0.627 0.922 0.913 0.926 0.943 0.803
SuGaR 0.639 0.486 0.776 0.686 0.566 0.910 0.892 0.908 0.932 0.755
Ours 0.731 0.573 0.845 0.764 0.630 0.918 0.908 0.927 0.940 0.804
3DGS 0.265 0.377 0.147 0.266 0.362 0.231 0.212 0.138 0.214 0.246
SuGaR 0.344 0.416 0.220 0.335 0.429 0.245 0.232 0.164 0.221 0.290
Ours 0.271 0.378 0.138 0.263 0.369 0.214 0.197 0.125 0.194 0.239
Table 7. PSNR scores for Synthetic NeRF dataset (Mildenhall et al., 2021).
Mic Chair Ship Materials Lego Drums Ficus Hotdog Avg.
@@
2DGS (Ours) 35.09 35.05 30.60 29.74 35.10 26.05 35.57 37.36 33.07

### v2

@@ 6. Experiments
We now present evaluations of our 2D Gaussian Splatting reconstruction method, including appearance and geometry comparison with previous state-of-the-art implicit and explicit approaches. We then analyze the contribution of the proposed components.
@@ 6.1. Implementation
We implement our 2D Gaussian Splatting with custom CUDA kernels, building upon the framework of 3DGS (Kerbl et al., 2023). We extend the renderer to output depth distortion maps, depth maps and normal maps for regularizations (See detailed computations in Appendices A and B of the supplemental material). During training, we increase the number of 2D Gaussian primitives following the adaptive control strategy in 3DGS. Since our method does not directly rely on the gradient of the projected 2D center, we hence project the gradient of 3D center $\mathbf{p}_{k}$ onto the screen space as an approximation. Similarly, we employ a gradient threshold of $0.0002$ and remove splats with opacity lower than $0.05$ every $3000$ step. We conduct all the experiments on a single GTX RTX3090 GPU.
Mesh Extraction
To extract meshes from reconstructed 2D splats, we render depth maps of the training views using the depth value of the splats projected to the pixels and utilize truncated signed distance fusion (TSDF) to fuse the reconstruction depth maps, using Open3D (Zhou et al., 2018). We set the voxel size to $0.004$ and the truncated threshold to 0.02 during TSDF fusion. We also extend the original 3DGS to render depth and employ the same technique for surface reconstruction for a fair comparison.
@@
Courthouse 0.17 0.12 0.28 0.08 0.09 0.16
Ignatius 0.83 0.72 0.89 0.33 0.04 0.51
Meetingroom 0.24 0.20 0.32 0.15 0.01 0.17
Truck 0.45 0.45 0.48 0.26 0.19 0.45
Mean 0.38 0.35 0.50 0.19 0.09 0.32
Time ¿24h ¿24h ¿24h ¿1h 14.3 m 15.5  m
Table 3. Performance comparison between 2DGS (ours), 3DGS and SuGaR on the DTU dataset (Jensen et al., 2014). We report the averaged chamfer distance, PSNR (training-set view), reconstruction time, and model size.
@@
SuGaR 22.93 0.629 0.356 29.43 0.906 0.225
3DGS 24.64 0.731 0.234 30.41 0.920 0.189
2DGS (Ours) 24.34 0.717 0.246 30.40 0.916 0.195
6.2. Comparison
Dataset
We evaluate the performance of our method on various datasets, including DTU (Jensen et al., 2014), Tanks and Temples (Knapitsch et al., 2017), and Mip-NeRF360 (Barron et al., 2022a). The DTU dataset comprises 15 scenes, each with 49 or 69 images of resolution $1600\times 1200$. We use Colmap (Schönberger and Frahm, 2016) to generate a sparse point cloud for each scene and down-sample the images into a resolution of $800\times 600$ for efficiency. We use the same training process for 3DGS (Kerbl et al., 2023) and SuGaR (Guédon and Lepetit, 2023) for a comparison.
Geometry Reconstruction
In  Table 1 and  Table 3, we compare our geometry reconstruction to SOTA implicit (i.e., NeRF (Mildenhall et al., 2020), VolSDF (Yariv et al., 2021), and NeuS (Wang et al., 2021)), explicit (i.e., 3DGS (Kerbl et al., 2023) and concurrent work SuGaR (Guédon and Lepetit, 2023)) methods on Chamfer distance and training time using the DTU dataset. Our method outperforms all compared methods in terms of Chamfer distance. Moreover, as shown in Table 2, 2DGS achieves competitive results with SDF models (i.e., NeuS (Wang et al., 2021) and Geo-Neus (Fu et al., 2022)) on the TnT dataset, and significantly better reconstruction than explicit reconstruction methods (i.e., 3DGS and SuGaR). Notably, our model demonstrates exceptional efficiency, offering a reconstruction speed that is approximately 100 times faster compared to implicit reconstruction methods and more than 3 times faster than the concurrent work SuGaR. Our approach can also achieve qualitatively better reconstructions with more appearance and geometry details and fewer outliers, as shown in  Figure 5. Moreover, SDF-based reconstruction methods require predefining the spherical size for initialization, which plays a critical role in the success of SDF reconstruction. By contrast, our method leverages radiance field based geometry modeling and is less sensitive to initialization. We include the full geometry reconstruction results for both DTU and TnT in Figure 9 and Figure 10.
Appearance Reconstruction
Our method represents 3D scenes as radiance fields, providing high-quality novel view synthesis. In this section, we compare our novel view renderings using the Mip-NeRF360 dataset against baseline approaches, as shown in  Table 4 and  Figure 4. Note that, since the ground truth geometry is not available in the Mip-NeRF360 dataset and we hence focus on quantitative comparison. Remarkably, our method consistently achieves competitive NVS results across state-of-the-art techniques while providing geometrically accurate surface reconstruction. We include the appearance rendering results in Figure 11.
@@ 6.2. Comparison
Accuracy $\downarrow$ Completion  $\downarrow$ Average $\downarrow$
@@
6.3. Ablations
[…]
Regularization
We first examine the effects of the proposed normal consistency and depth distortion regularization terms. Our model (Table 5 E) provides the best performance when applying both regularization terms. We observe that disabling the normal consistency (Table 5 A) can lead to incorrect orientation, as shown in Figure 6 A. Additionally, the absence of depth distortion (Table 5 B) results in a noisy surface, as shown in Figure 6 B.
Mesh Extraction
We now analyze our choice for mesh extraction. Our full model (Table 5 E) utilizes TSDF fusion for mesh extraction with median depth. One alternative option is to use the expected depth instead of the median depth. However, it yields worse reconstructions as it is more sensitive to outliers, as shown in Table 5 C. Further, our approach surpasses screened Poisson surface reconstruction (SPSR)(Table 5 D) (Kazhdan and Hoppe, 2013) using 2D Gaussians’ center and normal as inputs, due to SPSR’s inability to incorporate the opacity and the size of 2D Gaussian primitives.
@@ 7. Conclusion
Limitations
While our method successfully delivers accurate appearance and geometry reconstruction for a wide range of objects and scenes, we also discuss its limitations: First, we assume surfaces with full opacity and extract meshes from multi-view depth maps. This can pose challenges in accurately handling semi-transparent surfaces, such as glass, due to their complex light transmission properties, as shown in Figure 12. Secondly, our current densification strategy favors texture-rich over geometry-rich areas, occasionally leading to less accurate representations of fine geometric structures. A more effective densification strategy could mitigate this issue. Finally, our regularization often involves a trade-off between image quality and geometry, and can potentially lead to over-smoothing in certain regions.
Acknowledgements.
BH and SG are supported by NSFC #62172279, #61932020, Program of Shanghai Academic Research Leader. ZY, AC and AG are supported by the ERC Starting Grant LEGO-3D (850533) and DFG EXC number 2064/1 - project number 390727645.
@@ Appendix A Details of Depth Distortion
(17) $\begin{split}\mathcal{L}&=\sum_{i=0}^{N-1}\sum_{j=0}^{i-1}\omega_{i}\omega_{j}(m_{i}-m_{j})^{2}\\ &=\sum_{i=0}^{N-1}\omega_{i}\left(m_{i}^{2}\sum_{j=0}^{i-1}\omega_{j}+\sum_{j=0}^{i-1}\omega_{j}m_{j}^{2}-2m_{i}\sum_{j=0}^{i-1}\omega_{j}m_{j}\right)\\ &=\sum_{i=0}^{N-1}\omega_{i}\left(m_{i}^{2}A_{i-1}+D^{2}_{i-1}-2m_{i}D_{i-1}\right),\end{split}$
@@
Figure 7. Visualization of a plane tiled by 2D Gaussians. Affine approximation (Zwicker et al., 2001b) adopted in 3DGS (Kerbl et al., 2023) causes perspective distortion and inaccurate depth, violating normal consistency.
@@ Appendix B Depth calculations
Mean depth:
There are two optional depth computations used for our meshing process. The mean (expected) depth is calculated by weighting the intersected depth:
@@
Median depth:
We compute the median depth as the largest “visible” depth, considering $T_{i}=0.5$ as the pivot for surface and free space:
@@
(a) Ground-truth
(b) MipNeRF360 (Barron et al., 2022b), SSIM=0.813
  (c) 3DGS, normals from depth gradientRefer to caption Refer to caption (d) 3DGS (Kerbl et al., 2023), SSIM=0.834   (e) Our model (2DGS), normals from depth gradientRefer to caption Refer to caption (f) Our model (2DGS), SSIM=0.845
Figure 8. We visualize the depth maps generated by MipNeRF360 (Barron et al., 2022b), 3DGS (Kerbl et al., 2023), and our method. The depth maps for 3DGS (d) and 2DGS (f) are rendered using Eq. 18 and visualized following MipNeRF360. To highlight the surface smoothness, we further visualize the normal estimated from depth gradient using Eq. 15 for both 3DGS (c) and ours (e). While MipNeRF360 is capable of producing plausibly smooth depth maps, its sampling process may result in the loss of detailed structures. Both 3DGS and 2DGS excel at modeling thin structures; however, as illustrated in (c) and (e), the depth map of 3DGS exhibits significant noise. In contrast, our approach generates sampled depth points with normals consistent with the rendered normal map (refer to Figure 1b), thereby enhancing depth fusion during the meshing process.
Appendix C Additional baselines
In this section, we present additional baselines to ablate the impact of our design choices, as summarized in Table 6. Furthermore, we integrate our meshing approach into the comparison against these baselines for a comprehensive analysis. SuGaR extracts a mesh from depth points utilizing SPSR (Screen Poisson Surface Reconstruction) during the coarse stage, followed by refinement using a mesh renderer. To assess the effect of this meshing strategy, we substituted SPSR with TSDF using the depth maps, followed by an identical refinement stage. However, we found that the depth map generated from their flat Gaussian intersection is sparse and discontinuous. As a result, the adaptation of TSDF with its discontinuous depth map yields inferior results. For 3DGS, we leverage SPSR for mesh generation. Because the 3D Gaussian lacks a surface normal, we treat its normal as a trainable parameter (Gao et al., 2023; Liang et al., 2023) distilled from the depth map, employing the normal consistency regularization. We then utilize all center points for SPSR, resulting in improved overall completion metrics compared to results obtained using TSDF.
Finally, we conduct ablation experiments on our 2DGS. Notably, 2DGS demonstrates enhanced performance by iteratively integrating components such as TSDF, perspective-correct rasterization, and median depth. For the affine approximation baseline, we utilize 3DGS’s rasterization method by configuring one scale of the 3D Gaussian to $1e^{-6}$. While affine approximation already yields promising results, integrating the proposed ray-splat intersection scheme results in more accurate depth map generation under perspective projection, as depicted in Figure 7, thus enhancing depth fusion performance.
@@ Appendix C Additional baselines
Accuracy $\downarrow$ Completion $\downarrow$ Average $\downarrow$
SuGaR 1.48 1.17 1.33
SuGaR + TSDF 2.47 1.90 2.18
3DGS + SPSR (center) 2.05 1.25 1.65
3DGS + TSDF (mean) 1.93 1.99 1.96
2DGS + SPSR (center) 1.25 0.89 1.07
2DGS (affine approx) + TSDF (mean) 0.96 1.20 1.08
2DGS (our rasterizer) + TSDF (mean) 0.79 0.98 0.88
2DGS (our rasterizer) + TSDF (median) 0.78 0.83 0.80
Appendix D Additional Results
Our 2D Gaussian Splatting method achieves comparable performance even without the need for regularizations, as Table 7 shows. We have included a detailed breakdown of per-scene metrics for the MipNeRF360 dataset (Barron et al., 2022b) in Table 9. Additionally, we have provided a comparison of our rendered depth maps with those from 3DGS and MipNeRF360 in Figure 8.
Table 7. PSNR scores for Synthetic NeRF dataset. Our model achieve comparable performance without using regularizations.
Mic Chair Ship Materials Lego Drums Ficus Hotdog Mean
@@ Appendix D Additional Results
Ours 35.09 35.05 30.60 29.74 35.10 26.05 35.57 37.36 33.07
Table 8. PSNR scores for TnT dataset.
Barn Caterpillar Courthouse Ignatius Meetingroom Truck Mean
SuGaR 28.63 23.27 23.33 20.72 25.47 24.40 24.16
3DGS 27.99 24.82 23.33 23.95 26.89 25.01 25.33
Ours 28.79 24.23 23.51 23.82 26.15 26.85 25.56
[…]
SugaR 23.34 19.54 25.40 25.07 21.30 29.97 27.56 29.41 30.77 25.82
3DGS 25.24 21.52 27.41 26.55 22.49 30.63 28.70 30.32 31.98 27.20
Ours 24.87 21.15 26.95 26.47 22.27 31.06 28.55 30.50 31.52 27.03
SuGaR 0.634 0.499 0.762 0.705 0.546 0.904 0.885 0.902 0.933 0.752
3DGS 0.771 0.605 0.868 0.775 0.638 0.914 0.905 0.922 0.938 0.815
Ours 0.752 0.588 0.852 0.765 0.627 0.912 0.900 0.919 0.933 0.805
SuGaR 0.354 0.407 0.240 0.325 0.452 0.259 0.244 0.178 0.220 0.298
3DGS 0.205 0.336 0.103 0.210 0.317 0.220 0.204 0.129 0.205 0.214
Ours 0.218 0.346 0.115 0.222 0.329 0.223 0.208 0.133 0.214 0.223